# MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

### v2

@@ 3.1 Mimicking 3D reconstruction to reduce the NFEs
For simplicity, let $x^{(k)}$ be a vector by flattening all images and concatenating them together and so do any subsequent set of multi-view images.
As detailed in Sec. 3.2, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [42, 6, 44, 43]. The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for reconstruction in the current iteration. In particular, starting from $\theta_{0}^{(k)}$, keeping the camera poses $c^{(k)}$ unchanged, we optimize the 3D parameters w.r.t. the following reconstruction loss for $I^{(k)}$ steps:
@@
Here we assume the rendered images and those sampled by the diffusion model are the same dimensional for simplicity. In Appendix B.1 we will further discuss how to deal with the latent diffusion model (LDM) [17]. Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ for reconstruction in Sec. 3.2.
3.2 Refining multi-view images for reconstruction by diffusion
We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process (see Fig. 3b).
Figure 3: Overview of DDIM reconstruction and our refinement. (a) The inversion process can preserve the 3D consistency of input images, and the input images can be recovered through the corresponding DDIM sampler. (b) After the carefully designed refinement of $x$ by Eq. (9) and Eq. (10), the resulting $\hat{x}$ possesses enough high quality for 3D reconstruction.
The diffusion model has two equivalent forward processes [16]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq. (1) and the inversion process based on the probability flow ordinary differential equations (ODEs) in Eq. (4). In comparison, the noise-adding process is more efficient without function evaluation, and it aids in eliminating low-quality details through high-frequency noise. Conversely, the inversion process can preserve the 3D consistency and overall information of the current 3D object (see Fig. 3a). In addition, observations from existing experiments  [55] indicate that diffusion models tend to generate details closer to time 00 and overall structure as they approach time $T$.
Inspired by this, we design a hybrid forward process that initially adds noise to time $t_{1}^{(k)}$ and then perform DDIM inversion to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise.. In contrast to random sampled $t_{2}^{(k)}$ in existing algorithms including SDS [1], we adopt a linearly decreased schedule for $t_{2}^{(k)}$ as $k$ progresses, detailed in Sec. 4. Formally, the process is defined as:
@@ 3.2 Refining multi-view images for reconstruction by diffusion
Note that generally $x^{(k)}$ does not follow the model distribution defined by the diffusion and the resulting $\tilde{x}^{(k)}$ do not strictly adhere to the marginal distribution at the corresponding time. Nevertheless, we still use existing sampling algorithms to obtain refined images $\hat{x}^{(k)}$ from $\tilde{x}^{(k)}$ as follows:
@@
Although other advanced sampling methods [46, 47, 48, 49, 50, 51] exist, we choose the popular DDIM in Eq. (3) and tune its noise hyperparameter $\eta$. The optimal value of $\eta$ depends on the base model and we search for the best one in $\{0,0.5,1\}$ as detailed in Tab. 1 in Appendix A.
@@
The whole process of SIR is presented in Algorithm 1. Compared with existing generation methods such as SDS, SIR benefits from reconstruction, i.e. leverages several NFEs for multiple updates of 3D parameters, and lowers the NFEs in total (see Fig. 4).
1
2Input: The number of iterations $K$, an initial 3D object $\theta_{0}^{(0)}$, a set of reconstruction steps $\{I^{(k)}\}_{k=0}^{K-1}$ and a set of camera poses $\{c^{(k)}\}_{k=0}^{K-1}$.
3 Output: A final 3D content $\theta^{(K)}_{0}$. for $k$ from 00 to $K-1$ do
4       Render $N$ images $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)})$ ;
5       Obtain noisy $\tilde{x}^{(k)}$ from forward process for $x^{(k)}$ using Eq. (9);
6       Obtain refined $\hat{x}^{(k)}$ from sampling process for $\tilde{x}^{(k)}$ using Eq. (10);
7       for $i$ from 00 to $I^{(k)}-1$ do
8             Compute loss $L=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|$ using Eq. (8);
9             Compute the gradient $\nabla L$, update $\theta^{(k)}_{i}$ to $\theta^{(k)}_{i+1}$;
10
11       end for
12      Obtain $\theta^{(k+1)}_{0}=\theta^{(k)}_{I^{(k)}}$.
13 end for
Algorithm 1 Score-based iterative reconstruction
4 MicroDreamer
@@
In addition to our algorithm SIR, we carefully analyze and design other critical elements in the 3D training and diffusion. Our comprehensive system is called MicroDreamer to highlight its efficiency.
3D initialization and camera views.
@@
We utilize a direct reconstruction approach to initialize the 3D content. Specifically, we optimize the loss function in Eq. (8) by several steps (see Tab. 1 in Appendix A for detailed values) to update the 3D parameter, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion model. The camera views are uniformly sampled following the corresponding baselines [9, 56] except that the azimuth angles of different views in the same iteration are evenly distributed. In this setting, we are unable to directly utilize the diffusion model in [38, 39, 40] due to their fixed camera views conditioned generation.
Annealed time schedule.
@@
We utilize an annealed time schedule $\{t_{2}^{(k)}\}_{k=0}^{K-1}$ for the end of the forward process. Intuitively, as the quality of the 3D assets improves, the input $x$ in Eq. (7) becomes more proximate to the model distribution, thus requiring fewer steps to refine. This differs from SDS, which samples uniformly random $t_{2}$. Our preliminary experiments in Fig. 7(c) suggest that a linearly annealed schedule is sufficient. The endpoints of the schedule depend on the diffusion model, detailed in Tab. 1 in Appendix A. See Appendix C for the schedule of $t_{1}$.
Reference image loss for image-to-3D.
A reference image is available in image-to-3D, which is regarded as the ground truth front view of the 3D object on Stable Zero123 [6, 43] following DreamGaussian [9]. In this way, we add the reference loss in the same form as the reconstruction loss with Eq. (8) in each iteration of training. The weight set for the reference loss can be seen in Tab. 1.
3DGS settings and mesh refinement.
For simplicity, our approach on 3DGS largely adheres to the settings in DreamGaussian [9] unless specified. We incorporate a densify and prune procedure at every 50 updates during the initial 200 updates. We follow the mesh extraction method in DreamGaussian  [9] and employ a threshold value of $0.2$ in the marching cube algorithm [57, 9]. We utilize SIR to optimize the exported mesh texture in 3 iterations, with 15 steps per iteration, and we use a noise-adding process and DDIM with $\eta=0$ for simplicity.
5 Experiment
We present the results on NeRF, 3DGS, and ablation sequentially. For a fair comparison, we utilize the widely used CLIP similarity [45] for quantitative comparison following DreamGaussian [9]. We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0. For all methods, the corresponding images are rendered from NeRF in Sec. 5.1 and mesh in Sec. 5.2. We report the average CLIP similarities between these images and the reference image or text. See Appendix D for details of the checkpoint and codebase.
5.1 Results on NeRF
We apply SIR algorithm on NeRF [35, 36, 56] leveraging three multi-view diffusion model from MVDream [42], Stable Zero123 [43, 6], and ImageDream [44]. For each model, we selected 6 input prompts from a widely used codebase [56] for testing, calculated NFEs, and recorded the change of the average CLIP similarity during the generation process. Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig. 4. Besides, SIR can generate NeRF of high visual quality, as detailed shown in Appendix E, and the conclusion agrees with the results in Fig. 4.
(a) Comparison on MVDream
(b) Comparison on Stable Zero123
(c) Comparison on ImageDream
Figure 4: Comparison of SIR and SDS on NeRF. The change of the CLIP similarity in the generation process and final NFEs on different models. SIR achieves a 5-20 times acceleration and lowers NFEs.
5.2 Results on 3D Gaussian splatting
Qualitative comparisons.
We show the generated 3D meshes between the fastest optimization-based baseline DreamGaussian [9] and ours in Fig. 5. MicroDreamer halves the generation time of DreamGaussian while retaining competitive performance. We provide more visualization results of our methods in Appendix E.
Figure 5: Qualitative comparisons on mesh from 3DGS. MicroDreamer retains a comparable performance in terms of the visual quality of the mesh with higher efficiency.
Quantitative comparisons.
We compare MicroDreamer with seven competitive baselines including Point-E [24], Shap-E [23], One-2-3-45 [31], TriplanceGaussian [33], Wonder3D [39], LGM [30] and DreamGaussian [9]. We record the optimization time  [^3] The detailed time recording method is provided in Appendix D. for each model on a single NVIDIA A100 (80GB) GPU, and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works [31, 38, 9], as shown in Fig. 7. As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity. We provide the time record and detailed tables in Appendix F. Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves. See detailed limitations and future work in Sec. 7 and Appendix E.
Figure 6: Quantitative comparisons. Our MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of CLIP similarity. The zero-shot methods are highlighted in red.
[…]
(a) Iterative reconstruction.
(b) Camera views.
(c) Time schedule.
(d) Forward process.
Figure 7: Ablation study. (a) Iterative reconstruction is necessary. (b) More camera views benefit. (c) A linear schedule is sufficient. (d) A hybrid forward process can be effective.
5.3 Ablation study
Ablation on iterative reconstruction.
Fig. 7(a) shows the necessity of iterative reconstruction. The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise. Although increasing $K$ improves quality, we set $K\leq 30$ for efficiency.
Ablation on number of camera poses.
SIR necessitates more than 1 camera pose for reconstruction, as shown in Fig. 7(b). The performance increases with more camera poses. We choose $6$ on 3DGS with Stable Zero123 [43] model to balance efficiency and quality. See Tab. 1 for values in other settings.
Ablation on time schedule.
We compare our linear schedule, the random schedule employed in SDS [1], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$. As shown in Fig. 7(c), the schedules perform similarly. For simplicity and efficiency, we use the linear schedule by default.
Ablation on forward process.
Fig. 7(d) compares three forward processes on 3DGS with the Stable Zero123 [43] model. The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec. 3.2. See Tab. 1 in Appendix A for other settings.
6 Related work
Optimization-based 3D generation.
Built upon text-to-image diffusion models, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 58, 59, 60] usually generate 3D objects without additional 3D data. Among them, the most relevant line of work [1, 8, 10, 4, 2] proposes various distillation algorithms. Besides SDS [1], ProlificDreamer [8] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference. LucidDreamer [10] incorporates DDIM inversion to strengthen the forward process. Though promising, such methods suffer from inefficiency problems due to large NFEs. Distinct from existing algorithms, SIR mimics 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion. Our experiments demonstrate that SIR is 5-20 times faster than SDS [1] on NeRF and twice as fast as the most competitive baseline DreamGaussian [9] on meshes.
Feed-forward 3D generation.
Contrary to optimization-based methods, feed-forward methods [20, 21, 22, 23, 24, 25, 26, 27, 28, 61, 29, 31, 32, 33, 34, 62, 63, 64, 65] elect to use large-scale 3D datasets [66, 67] training to achieve a process that can directly generate 3D objects. Early methods are characterized by their speed but often produce lower-quality 3D structures with simple textures. Several recent methods, utilizing more data, have exhibited possibilities. Large reconstruction model (LRM) [68, 26] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image. Instant3D [69] further expands on LRM enabling it to accept multi-view input and reconstruct it into a NeRF. Remarkably, MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [45, 9].
Multi-view prediction based 3D generation.
There is also a line of work [38, 39, 40, 41] dedicated to enhancing the output of multi-view diffusion models by training on 3D datasets to reconstruct 3D objects using a single reconstruction process without iteration. These methods like Wonder3D[39] typically rely on 3D prior knowledge and are limited to specific 3D representations with a long reconstruction time. In contrast, thanks to the iterative process in SIR, MicroDreamer applies to various 3D representations and is more efficient (see Fig. 7 for a direct comparison).
7 Conclusion
We introduce SIR, an efficient and general algorithm combining both the strengths of 3D reconstruction and iterative optimization to reduce total NFEs in zero-shot 3D generation. SIR achieves a 5 to 20 times speed increase for NeRF generation compared to SDS. Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches with a competitive generation quality.
Limitations and broader impact.
There is potential for further improving MicroDreamer’s efficiency via employing consistency models [70, 71] or alternative sampling models that require fewer steps [72, 73]. Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion. Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves. Also, MicroDreamer can provide convenience for artists to create 3D assets. However, as a generative approach, our method may also be used for fabricating data and news, which warrants careful consideration.
8 Acknowledgement
Thanks Cheng Lu for the discussion when he was a PhD student at Tsinghua University.
Appendix A Key hyperparameters of MicroDreamer
@@
Table 1: Key hyperparameters of MicroDreamer on three base diffusion models. All models are employed to generate NeRF and the last two are employed to generate 3DGS and mesh. The hyperparameters are shared across the two settings by default. Otherwise, those for 3DGS are shown in brackets. $T$ is end time of diffusion.
Model Select MVDream [42] Stable Zero123 [6, 43] ImageDream [44]
@@
Forward process hybrid hybrid hybrid (random noise only)
Time schedule of $t_{2}$ $0.8T\rightarrow 0.5T$ $0.8T\rightarrow 0.2T$ $0.8T\rightarrow 0.6T$ ($0.8T\rightarrow 0.4T$)
@@
Reference color loss - 0.1 (0.3) 0
Reference opacity loss - 0.001 (0.01) 0
Appendix B The impact of latent diffusion model on 3D generation
Instead of applying the diffusion model directly in the pixel space to generate images, existing research [17] tends to use a VAE [75, 76] encoder to compress high-dimensional image data into a low-dimensional latent space first. Then a diffusion process is employed within this latent space to generate new images, which are finally decoded back by a VAE decoder to the pixel space to achieve high-quality images. We refer to this type of diffusion model as the latent diffusion model (LDM). The framework of LDM is adopted widely [6, 42, 44, 43], which leads to a significant impact on 3D generation.
B.1 SIR enabling optimization in the pixel space
Within the framework of LDM [17], the rendered images are mapped through an encoder $\mathcal{E}$ [75, 76] to a latent space, where the loss function is calculated. Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation. In such a case, the SIR loss in the latent space is given by:
An alternative approach maps the diffusion output back to the pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space. This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
However, we argue that this is not feasible for existing optimization methods [1, 2, 8] such as SDS. Even reparameterized as the data predicting form in Eq. (6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix B.2). In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space. The SIR loss in the pixel space is formalized as:
which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as the ablation results shown in Fig. 8. It is set as the default loss throughout the paper.

### v3

@@ IV-B Mimicking 3D reconstruction to reduce the NFEs
For simplicity, let $x^{(k)}$ be a vector by flattening all images and concatenating them together, and so do any subsequent set of multi-view images.
As detailed in Sec. IV-C, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [67, 6, 69, 68]. The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for reconstruction in the current iteration. In particular, starting from $\theta_{0}^{(k)}$, keeping the camera poses $c^{(k)}$ unchanged, we optimize the 3D parameters w.r.t. the following reconstruction loss for $I$ steps:
@@
Here we assume the rendered images and those sampled by the diffusion model have the same dimensional for simplicity. In Sec. IV-D we will further discuss how to deal with the latent diffusion model (LDM) [23]. Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ for reconstruction in Sec. IV-C.
Figure 5: The hybrid forward process is more efficient than DDIM inversion and generates better samples than noise-adding. We present the final results and sampling time on 20 iterations of SIR for three forward processes. Notably, the Noise-adding process may generate artifacts that contain unexpected elements compared to the input.
IV-C Refining multi-view images for reconstruction by diffusion
Algorithm 1 Score-based iterative reconstruction (SIR)
1:  Input: The number of iterations $K$, an initial 3D object $\theta_{0}^{(0)}$, the number of reconstruction steps $I$ and a set of camera poses $\{c^{(k)}\}_{k=0}^{K-1}$.
2:  Output: A final 3D content $\theta^{(K)}_{0}$.
3:  for $k$ from 00 to $K-1$ do
4:     Render $N$ images $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)})$
5:     Obtain noisy $\tilde{x}^{(k)}$ from forward process for $x^{(k)}$
6:     Obtain refined $\hat{x}^{(k)}$ from sampling process for $\tilde{x}^{(k)}$
7:     for $i$ from 00 to $I-1$ do
8:        Compute loss $L=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|$
9:        Compute the gradient $\nabla L$, update $\theta^{(k)}_{i}$ to $\theta^{(k)}_{i+1}$
10:     end for
11:  end for
We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process.
The diffusion model has two theoretically equivalent forward processes [22]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq. (1) and the inversion process based on the probability flow ordinary differential equations (ODEs) in Eq. (4). In comparison, the noise-adding process is more efficient without function evaluation, but sampling after the noise-adding process may produce unexpected artifacts (see Fig. 5). The inversion process can better preserve the 3D consistency and overall information of the current 3D object, but it necessitates more NFEs with low efficiency. We carefully design a hybrid forward process that initially adds noise and then performs DDIM inversion. Compared to common noise-adding and inversion processes, the hybrid forward process achieves a better balance in quality and efficiency, as Fig. 5 shows.
Specifically, the hybrid forward process adds noise to time $t_{1}^{(k)}$ first and then performs DDIM inversion [80] to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise.. In contrast to random sampled $t_{2}^{(k)}$ in existing algorithms including SDS [1], we adopt a linearly decreased schedule for $t_{2}^{(k)}$ as $k$ progresses, detailed in Sec. V. Formally, the process is defined as:
@@ IV-C Refining multi-view images for reconstruction by diffusion
Note that generally $x^{(k)}$ does not follow the model distribution defined by the diffusion and the resulting $\tilde{x}^{(k)}$ does not strictly adhere to the marginal distribution at the corresponding time. Nevertheless, we still use existing sampling algorithms to obtain refined images $\hat{x}^{(k)}$ from $\tilde{x}^{(k)}$ as follows:
@@
Although other advanced sampling methods [74, 75, 76, 77, 78, 79] exist, we choose the popular DDIM in Eq. (3) and tune its noise hyperparameter $\eta$. The optimal value of $\eta$ depends on the base model and we search for the best one in $\{0,0.5,1\}$ as detailed in Tab. I.
@@
The whole process of SIR is presented in Algorithm 1. Compared with existing methods [1, 7, 9, 15] dedicated to enhancing 3D generation quality, SIR benefits from reconstruction and lowers the NFEs in total, consequently improving generation efficiency. (see Fig. 7 for the comparison to SDS).
We visualize the optimization process of SIR in Fig. 6. With the increase in the number of iterations, the visual quality of the 3D content improves, validating the effectiveness of our iterative reconstruction.
Figure 6: Visualization of the optimization process in SIR. The visual quality of the 3D samples increases along with the iterations.
IV-D Enabling optimization in pixel space
Within the widely adopted framework of LDM [23], the rendered images are mapped through an encoder $\mathcal{E}$ [54, 55] to a latent space, where the loss function is calculated. Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation. In such a case, the SIR loss in latent space is given by:
An alternative approach maps the diffusion output back to pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space. This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
As discussed in Sec. IV-A, previous optimization-based methods like SDS fail to be effective when applied in pixel space. In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space. The SIR loss in pixel space is formalized as:
which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as the analysis results shown in Fig. 10. It is set as the default loss throughout the paper.
V MicroDreamer
@@
We provide a general configuration in the 3D training and diffusion for the SIR algorithm. The comprehensive system is called MicroDreamer to highlight its efficiency.
@@
3D initialization and camera views. We utilize a direct reconstruction approach to initialize the 3D content. Specifically, we optimize the loss function in Eq. (8) by several steps (see Tab. I for detailed values) to update the 3D parameters, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion models. The camera views are uniformly sampled following the corresponding baselines [8, 84] except the azimuth angles of different views in the same iteration are evenly distributed. In this setting, we are unable to directly utilize the diffusion model in [59, 60, 62, 63] due to their fixed camera views conditioned generation.
@@
Annealed time schedule. We utilize an annealed time schedule $\{t_{2}^{(k)}\}_{k=0}^{K-1}$ for the end of the forward process. Intuitively, as the quality of the 3D assets improves, the input $x$ in Eq. (7) becomes more proximate to the model distribution, thus requiring fewer steps to refine. This differs from SDS, which samples uniformly random $t_{2}$. Our preliminary experiments in Fig. 12(c) suggest that a linearly annealed schedule is sufficient. The endpoints of the schedule depend on the diffusion model, detailed in Tab. I.
@@
TABLE I: Key hyperparameters of MicroDreamer on three base diffusion models. All models are employed to generate NeRF and the last two are employed to generate 3DGS and mesh. By default, the hyperparameters are shared across the two 3D representations. Otherwise, those for 3DGS are shown in brackets. $T$ is the end time of diffusion.
Model Select MVDream [67] Stable Zero123 [68] ImageDream [69]
@@
Forward process hybrid hybrid hybrid
Time schedule of $t_{2}$ $0.8T\rightarrow 0.5T$ $0.8T\rightarrow 0.2T$ ($0.9T\rightarrow 0.2T$) $0.8T\rightarrow 0.6T$ ($0.8T\rightarrow 0.4T$)
Time schedule of $t_{1}$ $t_{1}=t^{2}_{2}/T$ $t_{1}=0.6t_{2}$ $t_{1}=t^{2}_{2}/T$ ($t_{1}=0.6t_{2}$)
@@
Ref. color loss - 0.1 (0.3) 0
Ref. opacity loss - 0.001 (0.01) 0
Reference image loss for image-to-3D. A reference image is available in image-to-3D, which is regarded as the ground truth front view of the 3D object on Stable Zero123 [68] following DreamGaussian [8]. In this way, we add the reference loss in the same form as the reconstruction loss with Eq. (8) in each training iteration. The weight set for the reference loss can be seen in Tab. I.
3DGS settings and mesh refinement. For simplicity, our approach on 3DGS largely adheres to the settings in DreamGaussian [8] unless specified. We incorporate a densify and prune procedure at every 100 updates during the initial 300 updates. At the end of optimization, we do a last prune and remove potential white Gaussians with scales larger than 0.01. We follow the mesh extraction method in LGM [43] and employ a threshold value of $5$ in the marching cube algorithm [86]. We utilize SIR to optimize the exported mesh texture for one iteration with 30 steps, and we use a noise-adding process and DDIM with $\eta=0$ for simplicity.
VI Experiments
We present the experimental details and results on NeRF, 3DGS, and ablation sequentially.
VI-A Experimental details
We present key hyperparameters of MicroDreamer in Tab. I. For implementing SIR and SDS on NeRF, we choose the popularly used framework threestudio [84]. For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [8]. Most hyperparameters are not sensitive to the results, and we follow the previous framework [84, 8] for these settings. For important hyperparameters, including the number of iterations, camera views, and the time schedule, see Sec. VI-D for ablation. See Tab. II for details about the URL of the codebases and checkpoints we used in this paper.
For a fair comparison, we utilize the widely used CLIP similarity [71] for quantitative comparison following DreamGaussian [8]. We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0. For all methods, the corresponding images are rendered from NeRF in Sec. VI-B and mesh in Sec. VI-C. We report the average CLIP similarities between these images and the reference image or text.
TABLE II: Codebases and checkpoints. We provide URLs for the open-source assets we used in this paper.
Model URL
Codebases
[84] https://github.com/threestudio-project/threestudio
[8] https://github.com/dreamgaussian/dreamgaussian
Checkpoints
[87] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k
[67] https://github.com/bytedance/MVDream-threestudio
[68] https://huggingface.co/stabilityai/stable-zero123
[69] https://github.com/bytedance/ImageDream
Baselines
[30] https://github.com/openai/point-e
[29] https://github.com/openai/shap-e
[39] https://github.com/One-2-3-45/One-2-3-45
[37] https://github.com/VAST-AI-Research/TriplaneGaussian
[60] https://github.com/xxlong0/Wonder3D
[43] https://github.com/3DTopia/LGM
[70] https://github.com/3DTopia/OpenLRM
(a) Comparion on MVDream [67]
(b) Comparion on Stable Zero123 [68]
(c) Comparion on ImageDream [69]
Figure 7: Comparison of SIR and SDS on NeRF. We plot the curve of the CLIP similarity in the generation process and final NFEs on different models. SIR lowers the NFEs and achieves a 5-20 times acceleration to achieve a competitive quality.
Figure 8: Qualitative comparison on NeRF. SIR can generate NeRF of higher visual quality than SDS in a short time.
VI-B Results on NeRF
We apply SIR algorithm on NeRF [57, 58, 84] leveraging three multi-view diffusion models from MVDream [67], Stable Zero123 [68], and ImageDream [69]. For each model, we selected 6 input prompts from a widely used codebase [84] for testing, calculated NFEs, and recorded the average CLIP similarity during the generation process. Compared with SDS, SIR lowers the total NFEs and accelerates the generation process 5-20 times while holding a competitive CLIP similarity, as shown in Fig. 7.
To provide a comprehensive comparison, we qualitatively analyze the generation results of SDS and SIR under identical time conditions, as illustrated in Fig. 8. Notably, at the moment of convergence in SIR, we find that SDS has not achieved satisfactory results, such as producing ambiguous outcomes. These observations support the quantitative results, demonstrating that SIR is general and more efficient.
VI-C Results on 3D Gaussian splatting
Qualitative comparisons. In Fig. 9, we present the generated 3D meshes comparing the fastest optimization-based baseline, DreamGaussian [8], with our method. MicroDreamer reduces DreamGaussian’s generation time by half while exhibiting superior performance, including improved texture and enhanced geometric structure.
Figure 9: Qualitative comparisons on mesh from 3DGS. MicroDreamer produces superior meshes, characterized by enhanced texture and reduced geometric artifacts, with greater efficiency than DreamGaussian.
Quantitative comparisons. We compare MicroDreamer with eight competitive baselines including Point-E [30], Shap-E [29], One-2-3-45 [39], TriplaneGaussian [^3] As TriplaneGaussian has no official mesh export code, we apply the mesh exported code from LGM [43] for it. [37], Wonder3D [60], LGM [43], Open-LRM [70] and DreamGaussian [8]. We record the generation time for each model on a single NVIDIA A100 (80GB) GPU and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works [39, 59, 8], as shown in Tab. III. MicroDreamer generates significantly better 3D content than DreamGaussian, as indicated by the standard deviation across multiple runs, and has higher efficiency.
In addition, MicroDreamer is on par with speed compared to feed-forward methods [70] trained on a substantial amount of 3D data and has a very competitive CLIP similarity. In conclusion, all those results suggest that SIR is a promising approach for efficient 3D generation.
TABLE III: Quantitative comparisons. MicroDreamer significantly outperforms the strong optimization-based baseline DreamGaussian in quality and efficiency and remains competitive with feed-forward methods. All results are averaged over three runs.
Method CLIP similarity $\uparrow$ Generation time $\downarrow$
Point-E [30] $0.566\pm 0.0011$ $\sim 24$s
Shap-E [29] $0.626\pm 0.0030$ $\sim 5$s
One-2-3-45 [39] $0.617\pm 0.0025$ $\sim 42$s
Wonder3D [60] $0.696\pm 0.0017$ $\sim 170$s
TriplaneGaussian [37] $0.691\pm 0.0016$ $\sim 7$s
LGM [43] $0.700\pm 0.0017$ $\sim 3$s
Open-LRM-mix-base-1.1[70] $0.704\pm 0.0000$ $\sim 23$s
DreamGaussian [8] $0.692\pm 0.0015$ $\sim 30+10$s
DreamGaussian-300 iter [8] $0.641\pm 0.0032$ $\sim 18+10$s
MicroDreamer-20 iter (Ours) $0.697\pm 0.0009$ $\sim 18+3$s
MicroDreamer-30 iter (Ours) $0.711\pm 0.0007$ $\sim 26+3$s
VI-D Analysis and ablation study
(a) On Stable Zero123[68].
(b) On ImageDream [69].